OpenStudy (anonymous):

What is the value of f(2) for the function f(x) = 5x + 3?

Parth (parthkohli):

Think that you are putting x in a box and getting 5 times x and 3 added. Same way, you'll get 5 times 2 and 3 added. 5*2 + 3 = 10 + 3 = 13

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow f(2) = 13 }\)

OpenStudy (anonymous):

thank you